# Ground state properties of the Heisenberg-compass model on the square lattice

### v1

@@ V.1 Order by disorder from thermal fluctuations
To determine ObD from thermal fluctuations at low temperatures for a given $\xi$, we consider a classical spin-wave expansion about each of the ordered ground states for that $\xi$, and investigate how the spin-wave excitations contribute to the free energy of the system Kawamura (1984); McClarty et al. (2014); Seabra et al. (2016); Yan et al. (2017). We assume that we are at sufficiently low temperature such that there are only small fluctuations (spin-wave excitations) about a particular classical ground state, i.e. a harmonic expansion is valid. The accidentally degenerate ground state for which the free energy of the harmonic spin wave fluctuations is minimal is the one selected by ObD at low temperatures. This (classical) free energy can alternatively be obtained from the quantum non-interacting or linear spin-wave theory, as shown in Ref. Seabra et al. (2016). If, for a given $\xi$, the linear spin-wave spectrum at wave vector $\bm{q}$ about a classical ordered state parametrized by angle $\phi$ is $\omega_{\bm{q}}(\phi)$, the classical free energy is given by $\cramped{F(\phi)=T\sum_{\bm{q}}\ln\omega_{\bm{q}}(\phi)}$ Seabra et al. (2016) where $\omega_{\bm{q}}(\phi)$ implicitly depends on $\xi$. A quantum linear spin-wave analysis using the Holstein-Primakoff formalism Auerbach (1998) is discussed in Appendix B, deriving the frequencies of the linear spin-wave modes in the different parameter regimes. This alternative route to compute the free energy of the classical spin waves is convenient as we shall reuse the results of the quantum linear spin wave analysis in the context of ObD selection from quantum fluctuations in Sec. VI.1 and Sec. VI.2.
@@
In Regime-I, II, IV, and V, where there is an accidental $O(2)$ degeneracy among the classical ground states, the spin-wave free energy $F(\phi)$ depends on the ground state (parametrized by $\phi$) about which the spin-wave analysis is performed. Minima are found only about certain discrete states out of the continuous $O(2)$ manifold of states in all of the above four regimes, a demonstration of ObD selection. For any $\xi$ in Regime-I, the free energy is minimized over four Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$, corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (5). Note that these four states are related to one another by a $C_{4}$ rotation about the $\bm{\hat{z}}$ axis, a consequence of the exact $C_{4}$ symmetry of the Hamiltonian in Eq. (1). Now, applying the duality transformation [Eq. (2)] to the states selected by ObD in Regime-I maps them to stripe states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (6) [Stripe-$\parallel$], giving the corresponding ObD selected states for any $\xi$ in Regime-II. In Regime-III, the low-temperature classical phase is described by two discrete ground states – the ferromagnetic states along ±$\pm\bm{\hat{z}}$ directions. The Klein duality between Regime-III and VI demands that the phase in Regime-VI is given by two Néel states along ±$\pm\bm{\hat{z}}$ directions. Finally, in Regime-IV, the classical spin-wave free-energy, $F(\phi)$, is minimized and gives ObD selection for the four ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (7). By the Klein duality, in Regime-V, ObD selects four stripe states corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (8) [Stripe-$\perp$]. The above analysis determines the complete low-temperature classical phase diagram of the Heisenberg-compass model on the square lattice. A representative state of the symmetry related classical phases selected by ObD for each of the four parameter regimes with accidential degeneracy is illustrated beside each of the corresponding regimes in Fig. 2.
@@ V.2 Classical Monte Carlo
  • $\bullet$
    Regime-II: We define a different magnetization, motivated from Eq. (6), $\bm{m}_{\rm II}\equiv\frac{1}{N}\sum_{\bm{r}}\Big{(}(-1)^{m}S^{x}_{\bm{r}},(-1)^{n}S^{y}_{\bm{r}},0\Big{)}$ with $\bm{r}=m\,\bm{x}+n\,\bm{y}$ and $m$, $n$ take values $0,1,\cdots,(L-1)$ for an $L\times L$ square lattice. We then define the order parameter, $O_{\rm II}\equiv\sqrt{(m_{\rm II}^{x})^{2}+(m_{\rm II}^{y})^{2}}\cos(4\phi_{\rm II})$ where $\phi_{\rm II}\equiv\tan^{-1}(m_{\rm II}^{y}/m_{\rm II}^{x})$. Note that for ObD selected states in this regime, discussed in Sec. V.1, one has $O_{\rm II}=1$.
  • $\bullet$
@@
  • $\bullet$
    Regime-V: We define the order parameter in the same vein as we did for Regime-II. With the magnetization defined as $\bm{m}_{\rm V}\equiv\frac{1}{N}\sum_{\bm{r}}\Big{(}(-1)^{n}S^{x}_{\bm{r}},(-1)^{m}S^{y}_{\bm{r}},0\Big{)}$ with $\bm{r}=m\bm{\hat{x}}+n\bm{\hat{y}}$. The order parameter $O_{\rm V}\equiv\sqrt{(m_{\rm V}^{x})^{2}+(m_{\rm V}^{y})^{2}}\cos(4\phi_{\rm V})$ where $\phi_{\rm V}\equiv\tan^{-1}(m_{\rm V}^{y}/m_{\rm V}^{x})$, giving $O_{\rm V}=1$ for the ObD selected states in this regime, discussed in Sec. V.1.
  • $\bullet$
    Regime-VI: The order parameter is the $z$ component of the staggered magnetization, $O_{\rm VI}\equiv|(1/N)\sum_{\bm{r}}(-1)^{\bm{r}}S_{\bm{r}}^{z}|$, characterizing the Néel states along the ±$\pm\bm{\hat{z}}$ directions.
These order parameters along with the regime-relevant magnetizations are summarized in Table 1. Based on the low-temperature classical expansion described in Sec. V.1, we expect that the thermal average of the order parameters defined for a particular regime should approach unity in that regime (and approach zero elsewhere) as we move towards zero temperature. We measure the thermal averages of these order parameters using classical MC simulations at low temperatures. The simulations are performed on a square lattice with $N=L^{2}$ sites assuming periodic boundary conditions. The spins under consideration are the three component (Heisenberg) spins of unit length. The MC simulations are carried out based on an adaptive single-site Metropolis algorithm Alzate-Cardona et al. (2019), combined with over-relaxation moves Creutz (1987). We define a Monte Carlo sweep at a certain temperature as a combination of adaptive single-site Metropolis moves successively at $N$ randomly chosen sites with each followed by over-relaxation moves consecutively at 5 randomly chosen sites.
The full range of the phase angle $\xi\in[0,2\pi)$ is divided into 105 equally spaced points. For each value of $\xi$, we start with a random spin configuration at high temperature, $T=7$ and decrease to $T=2$ in temperature decrements of size $\delta T=0.1$, followed by a slower cooling down in steps of size $\delta T=0.01$ to a base temperature, $T=0.01$. In this way of cooling the system, at each temperature, we perform $5\times 10^{4}$ Monte Carlo sweeps for equilibration and then, measure the thermal averages of the above six order parameters as well as the specific heat, $C_{v}$  [^1] Specific heat at temperature $T$ from Monte Carlo simulations is computed as $C_{v}=\frac{\langle E^{2}\rangle-\langle E\rangle^{2}}{T^{2}}$ where $\langle E\rangle$ and $\langle E^{2}\rangle$ are Monte Carlo averages of energy and energy-square at $T$, respectively., over $10^{6}$ MC samples, skipping three MC sweeps in between consecutive measurements.
@@
We briefly mention here some of the key points known about the phase at the special parameters $\xi=0,\pi/2,\pi$, and their Klein duals, i.e., the six phase boundaries. In the thermodynamic limit, there is no long-range ordering at any non-zero temperature with $\xi=0,\pi$, or at their Klein duals, as the Mermin-Wagner-Hohenberg theorem forbids spontaneously broken continuous symmetry at nonzero temperature in two dimensions  Mermin and Wagner (1966); Hohenberg (1967). These disordered phases would be eliminated by any infinitesimally weak anisotropy, giving way to long-range ordered phases. At the compass points, a directional or nematic ordering along the director $\bm{\hat{x}}$ or $\bm{\hat{y}}$ is found at low temperatures ($\cramped{T<T_{c}}$), with a phase transition into a disordered paramagnetic phase at high temperatures ($\cramped{T>T_{c}}$ishra et al. (2004); Wenzel and Janke (2008)  [^3] This phase transition falls in the two-dimensional Ising model universality class Mishra et al. (2004); Wenzel and Janke (2008).. Ref. [Mishra et al., 2004] argues that the nematic ordering reduces to a conventional long-range ordering at low temperatures in the presence of a weak $XY$ exchange which favors long-range order along the ±$\pm\bm{\hat{x}}$ or ±$\pm\bm{\hat{y}}$ directions. In the context of the present study, the relevant non-compass perturbation would be the Heisenberg interaction. However, the nematic ordering should occupy a finite temperature fan ($\cramped{T_{c}>T>J}$) that extends away from the compass points.
@@
It is of interest to consider the critical temperature, $T_{c}$, for the transition between the high-temperature paramagnetic phase and the low-temperature ordered phase in the six regimes found from the MC simulations. For each $\xi$, this temperature is obtained from the location of the peak of the specific heat data. We present $T_{c}$ vs $\xi$ data for two different system sizes, $L=16$ and 20, in Fig. 3(d). The existence of a single peak in $C_{v}$ in the wide range of temperature shown in Fig. 3(d) and the smooth change of the order parameter with $T$ below the peak reveals the presence of a single phase transition from the paramagnetic phase to the ordered phase as temperature is decreased from above $T_{c}$ toward zero temperature. Thus, for each $\xi$, there is a single magnetic ordered phase below $T_{c}$, starting to develop at $T_{c}$ and gradually strengthening as $T\rightarrow 0$. In other words, there is no further phase transition at any temperature below $T_{c}$ to any other phases from that developed just below $T_{c}$. As mentioned earlier, this single magnetic ordered phase below $T_{c}$ is consistent with the ordering found from the low temperature expansion valid at $T\ll T_{c}$ described in Sec. V.1.
@@
It may seem surprising that the numerically observed critical temperatures at $\xi=0,\pi$, and at their Klein dual points are significantly different from zero [see Fig. 3(d)], where there should not be any finite temperature phase transition due to the Mermin-Wagner-Hohenberg theorem Mermin and Wagner (1966); Hohenberg (1967) [^4] Note that at $\xi=0,\pi$ and at their Klein dual points, the model [Eq. (1)] is an $O(3)$ Heisenberg model, and not an $XY$ model that would have a Berezinskii–Kosterlitz–Thouless-type transition. The apparent non-zero $T_{c}$ at those points is a finite-size effect, which should go to zero in the thermodynamic limit $\cramped{(N\rightarrow\infty)}$. However, eliminating these finite size effects would be quite computationally challenging Schmoll et al. (2021) due to the infrared fluctuations diverging logarithmically in $L$. Thus, showing the vanishing of the transition temperature at these isotropic points in the parameter space would require significantly larger system sizes. Elsewhere in the parameter space, the Hamiltonian is anisotropic and thus, Mermin-Wagner-Hohenberg theorem does not prohibit ordering at non-zero temperature  [^5] Note that in Regime-I, II, IV, and V, potential infrared divergence from the low dimensionality of the lattice and accidental continuous ground state degeneracy gets eliminated by order by thermal disorder induced anisotropy, and the long-range ordering prevails. See Ref. [Khatua et al., 2023] for a discussion on the subtle competition between infrared-divergent fluctuations and the effective anisotropy induced via order by thermal disorder.. Therefore, $T_{c}$ is expected to converge as the system size takes a moderately large value away from these fine-tuned isotropic limits. This can be seen in the simulation results; the almost overlapping $T_{c}$ data for $L=16$ and 20 in Fig. 3(d) suggests that $T_{c}$ has nearly converged even at the relatively small size of $L=16$ for all $\xi$ values except for $\xi=0,\pi$, and their Klein dual points where $T_{c}$ should vanish in the thermodynamic limit.

### v2

@@ V.1 Order by disorder from thermal fluctuations
To determine ObD from thermal fluctuations at low temperatures for a given $\xi$, we consider a classical spin-wave expansion about each of the ordered ground states for that $\xi$, and investigate how the spin-wave excitations contribute to the free energy of the system Kawamura (1984); McClarty et al. (2014); Seabra et al. (2016); Yan et al. (2017). We assume that we are at sufficiently low temperature such that there are only small fluctuations (spin-wave excitations) about a particular classical ground state, i.e. a harmonic expansion is valid. The accidentally degenerate ground state for which the free energy of the harmonic spin wave fluctuations is minimal is the one selected by ObD at low temperatures. This (classical) free energy can alternatively be obtained from the quantum non-interacting or linear spin-wave theory, as shown in Ref. Seabra et al. (2016). If, for a given $\xi$, the linear spin-wave spectrum at wave vector $\bm{q}$ about a classical ordered state parametrized by angle $\phi$ is $\omega_{\bm{q}}(\phi)$, the classical free energy is given by $\cramped{F(\phi)=T\sum_{\bm{q}}\ln\omega_{\bm{q}}(\phi)}$ McClarty et al. (2014); Seabra et al. (2016) where $\omega_{\bm{q}}(\phi)$ implicitly depends on $\xi$. A quantum linear spin-wave analysis using the Holstein-Primakoff formalism Auerbach (1998) is discussed in Appendix B, deriving the frequencies of the linear spin-wave modes in the different parameter regimes. This alternative route to compute the free energy of the classical spin waves is convenient as we shall reuse the results of the quantum linear spin wave analysis in the context of ObD selection from quantum fluctuations in Sec. VI.1 and Sec. VI.2.
@@
In Regime-I, II, IV, and V, where there is an accidental $O(2)$ degeneracy among the classical ground states, the spin-wave free energy $F(\phi)$ depends on the ground state (parametrized by $\phi$) about which the spin-wave analysis is performed. Minima are found only about certain discrete states out of the continuous $O(2)$ manifold of states in all of the above four regimes, a demonstration of ObD selection. For any $\xi$ in Regime-I, the free energy is minimized over four Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$, corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (5). Note that these four states are related to one another by a $C_{4}$ rotation about the $\bm{\hat{z}}$ axis, a consequence of the exact $C_{4}$ symmetry of the Hamiltonian in Eq. (1). Now, applying the duality transformation [Eq. (2)] to the states selected by ObD in Regime-I maps them to stripe states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (6) (Stripe-$\parallel$), giving the corresponding ObD selected states for any $\xi$ in Regime-II. In Regime-III, the low-temperature classical phase is described by two discrete ground states – the ferromagnetic states along ±$\pm\bm{\hat{z}}$ directions. The Klein duality between Regime-III and VI demands that the phase in Regime-VI is given by two Néel states along ±$\pm\bm{\hat{z}}$ directions. Finally, in Regime-IV, the classical spin-wave free-energy, $F(\phi)$, is minimized and gives ObD selection for the four ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ directions corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (7). By the Klein duality, in Regime-V, ObD selects four stripe states corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ in Eq. (8) (Stripe-$\perp$). The above analysis determines the complete low-temperature classical phase diagram of the Heisenberg-compass model on the square lattice. A representative state of the symmetry related classical phases selected by ObD for each of the four parameter regimes with accidential degeneracy is illustrated beside each of the corresponding regimes in Fig. 2.
@@ V.2 Classical Monte Carlo
  2. (ii)
    Regime-II: Here, we define a different magnetization, motivated from Eq. (6), $\bm{m}_{\rm II}\equiv\frac{1}{N}\sum_{\bm{r}}\Big{(}(-1)^{m}S^{x}_{\bm{r}},(-1)^{n}S^{y}_{\bm{r}},0\Big{)}$ with $\bm{r}=m\,\bm{x}+n\,\bm{y}$ and $m$, $n$ take values $0,1,\cdots,(L-1)$ for an $L\times L$ square lattice. We then define the order parameter, $O_{\rm II}\equiv\sqrt{(m_{\rm II}^{x})^{2}+(m_{\rm II}^{y})^{2}}\cos(4\phi_{\rm II})$ where $\phi_{\rm II}\equiv\tan^{-1}(m_{\rm II}^{y}/m_{\rm II}^{x})$. Note that for ObD selected states in this regime, discussed in Sec. V.1, one has $O_{\rm II}=1$.
  3. (iii)
@@
  5. (v)
    Regime-V: Here, we define the order parameter in the same vein as we did for Regime-II. With the magnetization defined as $\bm{m}_{\rm V}\equiv\frac{1}{N}\sum_{\bm{r}}\Big{(}(-1)^{n}S^{x}_{\bm{r}},(-1)^{m}S^{y}_{\bm{r}},0\Big{)}$ with $\bm{r}=m\bm{\hat{x}}+n\bm{\hat{y}}$. The order parameter $O_{\rm V}\equiv\sqrt{(m_{\rm V}^{x})^{2}+(m_{\rm V}^{y})^{2}}\cos(4\phi_{\rm V})$ where $\phi_{\rm V}\equiv\tan^{-1}(m_{\rm V}^{y}/m_{\rm V}^{x})$, giving $O_{\rm V}=1$ for the ObD selected states in this regime, discussed in Sec. V.1.
  6. (vi)
    Regime-VI: For this regime, the order parameter is the $z$ component of the staggered magnetization, $O_{\rm VI}\equiv|(1/N)\sum_{\bm{r}}(-1)^{\bm{r}}S_{\bm{r}}^{z}|$, characterizing the Néel states along the ±$\pm\bm{\hat{z}}$ directions.
These order parameters along with the regime-relevant magnetizations are summarized in Table 1. Based on the low-temperature classical expansion described in Sec. V.1, we expect that the thermal average of the order parameters defined for a particular regime should approach unity in that regime (and approach zero elsewhere) as temperature is decreased towards zero. We measure the thermal averages of these various order parameters using classical MC simulations at low temperatures. The simulations are performed on a square lattice with $N=L^{2}$ sites assuming periodic boundary conditions. The spins under consideration are the three component (Heisenberg) spins of unit length (i.e. $|{\bm{S}}_{\bm{r}}|=1$). The MC simulations are carried out based on an adaptive single-site Metropolis algorithm Alzate-Cardona et al. (2019), combined with over-relaxation moves Creutz (1987). We define a Monte Carlo sweep at a certain temperature as a combination of adaptive single-site Metropolis moves successively at $N$ randomly chosen sites with each followed by over-relaxation moves consecutively at 5 randomly chosen sites.
The full range of the phase angle $\xi\in[0,2\pi)$ is divided into 105 equally spaced points. For each value of $\xi$, we start with a random spin configuration at high temperature, $T=7$ and decrease to $T=2$ in temperature decrements of size $\delta T=0.1$, followed by a slower cooling down in steps of size $\delta T=0.01$ to a base temperature of $T=0.01$. In this way of cooling the system, at each temperature, we perform $5\times 10^{4}$ Monte Carlo sweeps for equilibration and then, measure the thermal averages of the above six order parameters as well as the specific heat, $C_{v}$  [^1] Specific heat at temperature $T$ from Monte Carlo simulations is computed as $C_{v}=\frac{\langle E^{2}\rangle-\langle E\rangle^{2}}{T^{2}}$ where $\langle E\rangle$ and $\langle E^{2}\rangle$ are Monte Carlo averages of energy and energy-square at $T$, respectively., over $10^{6}$ MC samples, skipping three MC sweeps in between consecutive measurements.
@@
We briefly mention here some of the key points known about the phase at the special parameters $\xi=0,\pi/2,\pi$, and their Klein duals, i.e., the six phase boundaries. In the thermodynamic limit, there is no long-range ordering at any non-zero temperature with $\xi=0,\pi$, or at their Klein duals, as the Mermin-Wagner-Hohenberg theorem forbids spontaneously broken $O(3)$ continuous symmetry at nonzero temperature in two dimensions  Mermin and Wagner (1966); Hohenberg (1967). These disordered phases would be eliminated by any infinitesimally weak anisotropy, giving way to long-range ordered phases. At the compass points, a directional or nematic ordering along the director $\bm{\hat{x}}$ or $\bm{\hat{y}}$ is found at low temperatures ($\cramped{T<T_{\textrm{dir}}}$), with a phase transition into a disordered paramagnetic phase at high temperatures ($\cramped{T>T_{\textrm{dir}}}$ishra et al. (2004); Wenzel and Janke (2008)  [^3] This phase transition falls in the two-dimensional Ising model universality class Mishra et al. (2004); Wenzel and Janke (2008).. Ref. [Mishra et al., 2004] argues that the nematic ordering reduces to a conventional long-range ordering at low temperatures in the presence of a weak $XY$ exchange which favors long-range order along the ±$\pm\bm{\hat{x}}$ or ±$\pm\bm{\hat{y}}$ directions. In the context of the present study, the relevant non-compass perturbation would be the Heisenberg interaction. However, the nematic ordering should occupy a finite temperature fan ($\cramped{T_{\textrm{dir}}>T>J}$) that extends away from the compass points.
@@
It is of interest to consider the critical temperature, $T_{c}$, for the transition between the high-temperature paramagnetic phase and the low-temperature ordered phase in the six regimes found from the MC simulations. For each $\xi$, this temperature is obtained from the location of the peak of the specific heat data. We present $T_{c}$ vs $\xi$ data for two different system sizes, $L=16$ and 20, in Fig. 3(d). The existence of a single peak in $C_{v}$ in the wide range of temperature shown in Fig. 3(d) and the smooth change of the order parameter with $T$ below the peak indicates the presence of a single phase transition from the paramagnetic phase to the ordered phase at $T_{c}$. Thus, for each $\xi$, there is a single magnetic ordered phase below $T_{c}$, starting to develop at $T_{c}$ and gradually strengthening as $T\rightarrow 0$, with no further phase transition at any temperature below $T_{c}$ to any other phases than the one that developed at $T_{c}$. In other words, there is no further phase transition at any temperature below $T_{c}$ to any other phases from that developed just below $T_{c}$. As mentioned earlier, this single magnetic ordered phase below $T_{c}$ is consistent with the ordering found from the low temperature expansion valid at $T\ll T_{c}$ described in Sec. V.1.
@@
It may seem surprising at first sight that the numerically observed critical temperatures at $\xi=0,\pi$, and at their Klein dual points are significantly different from zero [see Fig. 3(d)], where there should not be any finite temperature phase transition in accordance with the Mermin-Wagner-Hohenberg theorem Mermin and Wagner (1966); Hohenberg (1967) [^4] Note that at $\xi=0,\pi$ and at their Klein dual points, the model [Eq. (1)] is an $O(3)$ Heisenberg model, and not an $XY$ model that would have a Berezinskii–Kosterlitz–Thouless-type transition. The apparent non-zero $T_{c}$ at those points is a finite-size effect, which should go to zero in the thermodynamic limit $\cramped{(N\rightarrow\infty)}$. However, eliminating these finite size effects would be quite computationally challenging Schmoll et al. (2021) due to the infrared fluctuations diverging only logarithmically in $L$ Thus, showing the vanishing of the transition temperature at these isotropic points in the parameter space would require significantly larger system sizes than the ones considered here. For other $\xi$ values, the Hamiltonian is anisotropic and thus the Mermin-Wagner-Hohenberg theorem does not prohibit ordering at non-zero temperature  [^5] Note that in Regime-I, II, IV, and V, potential infrared divergence from the low dimensionality of the lattice and accidental continuous ground state degeneracy gets eliminated by order by thermal disorder induced anisotropy, and the long-range ordering prevails. See Ref. [Khatua et al., 2023] for a discussion on the subtle competition between infrared-divergent fluctuations and the effective anisotropy induced via order by thermal disorder.. Therefore, $T_{c}$ is expected to converge as the system size takes a moderately large value away from these fine-tuned $\xi=0$, HAF*, $\pi$, and HF* isotropic limits. This can be seen in the simulation results; the almost overlapping $T_{c}$ data for $L=16$ and 20 in Fig. 3(d) suggests that $T_{c}$ has nearly converged even at the relatively small size of $L=16$ for all $\xi$ values except for $\xi=0,\pi$, and their Klein dual points where $T_{c}$ should vanish in the thermodynamic limit.